# Automatic Paper Reviewing with Heterogeneous Graph Reasoning over LLM-Simulated Reviewer-Author Debates

### v1

@@ Introduction
Peer review is essential to scientific progress, ensuring the quality, validity, and originality of research (alberts2008reviewing). High-quality reviews help guide the research community toward impactful contributions. However, the recent surge in paper submissions across disciplines (drozdz2024peer) has placed increasing strain on the peer review process. This exponential growth imposes a significant burden on human reviewers, making it increasingly difficult to ensure timely and consistent evaluations (StelmakhSSD21). Moreover, peer review is inherently subjective. In some cases, reviewers may be careless, biased, or even malicious, thereby undermining the fairness and reliability of the decision-making process (StelmakhSSD21; fairness-peer-review). These challenges have motivated the development of automatic reviewing systems as a promising solution to alleviate reviewer workload and promote more objective evaluations at scale.
@@
Current LLM-based automatic reviewing methods generally fall into two main categories: (1) Prompt-based approaches, which leverage the in-context learning capabilities of LLMs to generate review content without parameter updates. These include direct prompting methods (e.g., AI-Scientist (ai-scientist), which uses GPT-4 with tailored review templates) and multi-agent collaboration frameworks, where multiple LLM instances simulate interactions between reviewers and authors (e.g., AgentReview (agent-review) and ReviewMT (reviewMT)). (2) Fine-tuned approaches, which train open-source LLMs (e.g., LLaMA (llama) or Qwen (qwen) series) on peer review data to align the model’s outputs with expert reviewing criteria (e.g., CycleReviewer (weng2025cycleresearcher) and DeepReview (zhu2025deepreviewimprovingllmbasedpaper)).
Despite recent progress in automatic paper reviewing, existing methods still face several notable challenges. Prompt-based approaches, which rely solely on LLMs and instructions, often generate superficial and shallow review content (ZhouC024). Studies have shown that LLMs tend to produce generally positive but low-discriminative evaluations, failing to capture the nuanced reasoning required in peer review (grapheval). Moreover, these methods are highly sensitive to prompt design (ErricaSSB25); small changes in prompt wording can lead to significantly different outputs, resulting in limited stability and robustness. On the other hand, fine-tuned approaches that adapt open-source LLMs using peer review data suffer from data scarcity and bias. Publicly available high-quality review datasets are limited in size and scope, which restricts the generalizability of trained models. Additionally, these methods typically produce only a single-perspective review, lacking the ability to model multi-reviewer interactions and argumentative dynamics inherent in real-world peer review. Furthermore, both categories remain susceptible to hallucinations, wherein LLMs generate factually incorrect or misleading content, compromising fairness and reliability (HuangYMZFWCPFQL25).
@@
To evaluate the effectiveness of ReViewGraph for automatic paper reviewing, we collect three benchmark datasets from OpenReview and compare our approach against seven strong baseline methods. Experimental results demonstrate that ReViewGraph consistently outperforms all baselines across these datasets. Notably, it achieves an average relative improvement of 15.73% over the second-best baselines. Overall, the main contributions of this work are:
  • We propose ReViewGraph, a novel framework for automatic paper reviewing that models reviewer–author interactions as heterogeneous graphs constructed from LLM-simulated multi-round debates.
  • We design a structured heterogeneous debate graph with semantically-typed nodes and edges to capture fine-grained argumentative relations across diverse review perspectives, and leverage a graph neural network to perform relational reasoning over this structure.
  • Extensive experiments on three datasets show that ReViewGraph consistently outperforms 7 strong baselines, achieving an average relative improvement of 15.73% over the second-best models.
@@
Early efforts in automatic paper reviewing primarily relied on manually curated review data and traditional neural classifiers for acceptance prediction. KangADZKHS18 introduced the PeerRead dataset and trained classifiers using manually labeled reviews. GhosalVEB19 incorporated sentiment features to improve prediction accuracy. To move beyond simple acceptance prediction, researchers began exploring automatic review generation. BartoliLMT16 proposed one of the earliest neural frameworks, trained on 48 papers from their lab, to generate review comments. Nagata19 generated sentence-level feedback on grammar errors to support academic writing. ReviewRobot (WangZHKJR20) constructed a knowledge graph from the input paper, predicted review scores, and selected templated comments based on both scores and supporting evidence. However, traditional neural models struggled with long and technical documents, lacking the capacity for deep semantic understanding, which ultimately hindered progress in this field.
@@
Recent advancements in LLMs have inspired a growing body of work on LLM-based automatic reviewing. Multi-agent collaboration frameworks aim to simulate the multi-role dynamics of real-world peer review. For example, ReviewMT (reviewMT) and AIScientist (ai-scientist) reframe the review process as a multi-round, long-context dialogue among multiple roles, i.e., reviewers, authors, and meta-reviewers. AgentReview (agent-review) further explores this direction by prompting LLMs to assume diverse reviewer personalities and decision strategies. Fine-tuned reviewer models enhance alignment with human review standards through supervised training. CycleReviewer (weng2025cycleresearcher) fine-tunes an open-source LLM on domain-specific review data, simulating multiple reviewers who assess the paper across different dimensions. DeepReview (zhu2025deepreviewimprovingllmbasedpaper) extends this line of work by modeling multi-step reviewer reasoning and training on generated rationales to produce more coherent, logically grounded reviews. Graph-based approaches such as GraphEval (grapheval) leverage LLM prompting to segment abstracts into discrete opinion sentences, which are then connected via similarity-based edges and processed through graph reasoning to predict acceptance decisions.
@@
Figure 1: The workflow of ReViewGraph. Reviewer1, Reviewer2, and Reviewer3 represent three regular reviewer agents. Nodes are typed by shape: circles denote paper topic nodes, pentagrams denote evaluation dimension nodes (e.g., Methodological Novelty), squares denote reviewer agents (with different colors distinguishing different agents), and triangles represent the author agent.
@@
To represent the complex reasoning and interaction dynamics of the peer review process, we construct a heterogeneous debate graph from simulated reviewer-author interactions. Following the standard formulation of heterogeneous information networks (hgt), we define the graph as $G=\{\mathcal{V},\mathcal{E},\mathcal{A},\mathcal{R}\}$, where $\mathcal{V}$ is the set of nodes, $\mathcal{E}$ is the set of edges, $\mathcal{A}$ is the set of node types, and $\mathcal{R}$ is the set of edge types. Each node $v\in\mathcal{V}$ and $e\in\mathcal{E}$ is associated with a type via node and edge type mapping functions: $\psi(v):V\rightarrow\mathcal{A}$, $\eta(e):E\rightarrow\mathcal{R}$. Each edge corresponds to a meta-relation in the form $\psi(s),\eta(e),\psi(t)>$, where $s$ and $t$ are source and target nodes.
@@
Following prior works (agent-review; zhu2025deepreviewimprovingllmbasedpaper; weng2025cycleresearcher; grapheval), we use the full paper as input and aim to predict the final review decision. The decision $s$ is selected from a predefined set of review outcomes: accept or reject.
@@
We compare our proposed framework against several representative baselines that leverage LLMs for automatic paper review: (1) ICL-based Method (BrownMRSKDNSSAA20): Directly prompts a pre-trained LLM using in-context examples to predict the review decision based on the full paper content. (2) CoT-based Method (Wei0SBIXCLZ22): Encourages the LLM to perform chain-of-thought reasoning before generating the final review outcome. (3) AI-Scientist (ai-scientist): Redefines the review process as a multi-turn, long-context dialogue among multiple roles. (4) CycleReviewer (weng2025cycleresearcher): Fine-tunes an open-source LLM using domain-specific peer review data. (5) DeepReview (zhu2025deepreviewimprovingllmbasedpaper): Extends CycleReviewer by modeling multi-step reviewer reasoning and training on generated rationales. (6) GraphEval (grapheval): Segments paper abstracts into discrete opinion sentences using LLM prompting, connects them via similarity-based edges to construct a graph, and applies graph-based reasoning to predict the final review decision. The implementation details of ReViewGraph and the baselines are provided in Appendix.
@@
Case 1: Correct Rejection Prediction Amid Subtle Negative Consensus.
In the first case shown in Figure 2(a), the ground-truth decision was Reject. While the paper proposed a novel approach, multiple reviewers expressed concerns about the lack of rigorous theoretical analysis and inadequate comparisons with state-of-the-art baselines. Reviewer 1 noted the absence of evidence demonstrating the solution’s diversity. Reviewer 2 emphasized the theoretical gap in stochastic settings and the lack of comparison with methods designed for the $n\gg m$ scenario. Reviewer 3 echoed similar concerns regarding limited convergence analysis and experimental comparisons. Although the authors acknowledged these issues and promised improvements, no concrete resolutions were provided. ReViewGraph successfully captured the negative consensus across multiple evaluation dimensions (e.g., theoretical rigor, baseline comparison) by modeling reviewer–reviewer agreement and reviewer–author interactions in a structured heterogeneous graph. The system correctly identified the dominant negative stance despite superficially polite language or minor positive comments, and inferred the rejection outcome. In contrast, methods such as ICL, CoT prompting, and LLM-based dialogue agents misclassified the case as Accept, likely due to over-reliance on isolated affirmative signals.
Case 2: Correct Acceptance Prediction Despite Isolated Criticism.
In another instance shown in Figure 2(b), the ground-truth decision was Accept. Reviewer 2 suggested further exploration of the impact of calibration techniques on computational efficiency. However, both Reviewer 1 and Reviewer 3 explicitly stated that the paper adequately addressed computational concerns and praised its theoretical guarantees. Reviewer 2 ultimately agreed with the decision to accept. The authors also acknowledged the suggestion and expressed willingness to expand the analysis in future revisions. While several baseline models incorrectly predicted a Reject decision, possibly misinterpreting Reviewer 2’s suggestion as a critical flaw, our model accurately inferred that Reviewer 2’s comment was an isolated and non-decisive concern. By modeling the majority–minority stance across reviewers and the consistency between evaluation dimensions and final decisions, our method avoided overweighing minority dissent.
These case studies show that our approach captures structured reviewer interactions and evaluation alignment, enabling more robust and context-aware predictions.
@@ Conclusion
We proposed ReViewGraph, a novel framework for automatic paper reviewing that leveraged heterogeneous graphs constructed from simulated reviewer–author interactions. By modeling multi-agent debates, we extracted structured opinion relations among reviewers and authors and represented them as a heterogeneous graph with semantically typed nodes and edges. A heterogeneous graph Transformer was then applied to reason over this structure and predict final review outcomes. Through comprehensive experiments on three ICLR datasets, ReViewGraph consistently outperformed strong prompt-based, fine-tuned, and graph-based baselines across multiple evaluation metrics. Notably, our method achieved a relative improvement of 15.73% in Macro F1 score over the best-performing baseline, while requiring no LLM fine-tuning. Ablation studies further validated the importance of each graph component, particularly the modeling of evaluation dimensions and inter-agent relations. Case studies also demonstrated that ReViewGraph was able to correctly interpret nuanced reviewer dynamics, such as subtle consensus or isolated dissent, which were often misclassified by other methods.
These findings highlighted that explicitly modeling reviewer–author interactions and discourse-level semantics led to more robust, interpretable, and context-aware automated reviewing. Overall, ReViewGraph offered a scalable and effective framework for augmenting peer review with structured LLM reasoning, pointing toward future directions in trustworthy AI-assisted scientific evaluation.
@@ Acknowledgments
We would like to thank the anonymous reviewers for their valuable comments and suggestions to improve this paper. Min Yang was supported by National Key Research and Development Program of China (2024YFF0908200), National Natural Science Foundation of China (Grant No. 62376262), the Natural Science Foundation of Guangdong Province of China (2024A1515030166, 2025B1515020032). Xian Wei is supported by the General Program of Shanghai Natural Science Foundation (No.24ZR1419800, No.23ZR1419300), the National Natural Science Foundation of China (No.42130112, No.42371479), the Science and Technology Commission of Shanghai Municipality (No.22DZ2229004), and Shanghai Frontiers Science Center of Molecule Intelligent Syntheses. Shiwen Ni was supported by GuangDong Basic and Applied Basic Research Foundation (2023A1515110718 and 2024A1515012003), China Postdoctoral Science Foundation (2024M753398), Postdoctoral Fellowship Program of CPSF (GZC20232873) and Shenzhen Science and Technology Program (JCYJ20250604182917023).
@@
For structured reasoning over reviewer-author interactions, we design a heterogeneous graph and apply a two-layer Heterogeneous Graph Transformer (HGT) with a total of three layers. Each layer uses multi-head attention, and the hidden dimension is set to 128. To initialize node and edge representations, we use Qwen3-Embedding [^6] https://huggingface.co/Qwen/Qwen3-Embedding-8B. The model is trained for up to 100 epochs with a batch size of 32, using early stopping based on validation F1 score. Optimization is performed using the Adam optimizer (KingmaB14) with a learning rate of $1\times 10^{-4}$. Our code and data are provided in the supplementary materials.
@@
Table 9 and Table 10 respectively summarize the prompt construction and implementation details of the ICL-based (BrownMRSKDNSSAA20) and CoT-based methods (Wei0SBIXCLZ22). Table 11 and Table 12 present the prompt templates used by the AI-Scientist (ai-scientist) method for paper reviewing and review aggregation, respectively. All baselines are run using consistent inference settings (e.g., temperature, max tokens) unless otherwise stated.
For the GraphEval (grapheval) baseline, we follow the original implementation to preprocess our data and apply the same model configuration for training and evaluation [^7] https://github.com/ulab-uiuc/GraphEval. Specifically, we adopt a Graph Convolutional Network (GCN) as the backbone, configured as a two-layer weighted GNN with a hidden dimension of 64. The model is trained for up to 100 epochs with a batch size of 32, using the Adam optimizer.
For DeepReview (zhu2025deepreviewimprovingllmbasedpaper) and CycleReview (weng2025cycleresearcher), we directly invoke the publicly released models, DeepReviewer-14B [^8] https://huggingface.co/WestlakeNLP/DeepReviewer-14B, CycleReviewer-ML-LLaMA-3.1-8B [^9] https://huggingface.co/WestlakeNLP/CycleReviewer-ML-Llama-3.1-8B, and CycleReviewer-LLaMA-3.1-70B [^10] https://huggingface.co/WestlakeNLP/CycleReviewer-Llama-3.1-70B, without any fine-tuning or additional adaptation. This ensures a fair and reproducible baseline comparison, highlighting the benefits of our explicitly structured multi-agent debate framework.

### v2

@@ Introduction
Peer review is essential to scientific progress, ensuring the quality, validity, and originality of research (Alberts et al. 2008). High-quality reviews help guide the research community toward impactful contributions. However, the recent surge in paper submissions across disciplines (Drozdz and Ladomery 2024) has placed increasing strain on the peer review process. This exponential growth imposes a significant burden on human reviewers, making it increasingly difficult to ensure timely and consistent evaluations (Stelmakh et al. 2021). Moreover, peer review is inherently subjective. In some cases, reviewers may be careless, biased, or even malicious, thereby undermining the fairness and reliability of the decision-making process (Stelmakh et al. 2021; Zhang et al. 2022). These challenges have motivated the development of automatic reviewing systems as a promising solution to alleviate reviewer workload and promote more objective evaluations at scale.
@@
Current LLM-based automatic reviewing methods generally fall into two main categories: (1) Prompt-based approaches, which leverage the in-context learning capabilities of LLMs to generate review content without parameter updates. These include direct prompting methods (e.g., AI-Scientist (Lu et al. 2024), which uses GPT-4 with tailored review templates) and multi-agent collaboration frameworks, where multiple LLM instances simulate interactions between reviewers and authors (e.g., AgentReview (Jin et al. 2024) and ReviewMT (Tan et al. 2024)). (2) Fine-tuned approaches, which train open-source LLMs (e.g., LLaMA (Touvron et al. 2023) or Qwen (Bai et al. 2023) series) on peer review data to align the model’s outputs with expert reviewing criteria (e.g., CycleReviewer (Weng et al. 2025) and DeepReview (Zhu et al. 2025)).
Despite recent progress in automatic paper reviewing, existing methods still face several notable challenges. Prompt-based approaches, which rely solely on LLMs and instructions, often generate superficial and shallow review content (Zhou et al. 2024). Studies have shown that LLMs tend to produce generally positive but low-discriminative evaluations, failing to capture the nuanced reasoning required in peer review (Feng et al. 2025). Moreover, these methods are highly sensitive to prompt design (Errica et al. 2025); small changes in prompt wording can lead to significantly different outputs, resulting in limited stability and robustness. On the other hand, fine-tuned approaches that adapt open-source LLMs using peer review data suffer from data scarcity and bias. Publicly available high-quality review datasets are limited in size and scope, which restricts the generalizability of trained models. Additionally, these methods typically produce only a single-perspective review, lacking the ability to model multi-reviewer interactions and argumentative dynamics inherent in real-world peer review. Furthermore, both categories remain susceptible to hallucinations, wherein LLMs generate factually incorrect or misleading content, compromising fairness and reliability (Huang et al. 2025).
@@
To evaluate the effectiveness of ReViewGraph for automatic paper reviewing, we collect three benchmark datasets from OpenReview and compare our approach against seven strong baseline methods. Experimental results demonstrate that ReViewGraph consistently outperforms all baselines across these datasets. Notably, it achieves an average relative improvement of 15.73% over the second-best baselines. Overall, the main contributions of this work are: (1) We propose ReViewGraph, a novel framework for automatic paper reviewing that models reviewer–author interactions as heterogeneous graphs constructed from LLM-simulated multi-round debates. (2) We design a structured heterogeneous debate graph with semantically-typed nodes and edges to capture fine-grained argumentative relations across diverse review perspectives, and leverage a graph neural network to perform relational reasoning over this structure. (3) Extensive experiments on three datasets show that ReViewGraph consistently outperforms 7 strong baselines, achieving an average relative improvement of 15.73% over the second-best models.
@@
Early efforts in automatic paper reviewing primarily relied on manually curated review data and traditional neural classifiers for acceptance prediction. Kang et al. (2018) introduced the PeerRead dataset and trained classifiers using manually labeled reviews. Ghosal et al. (2019) incorporated sentiment features to improve prediction accuracy. To move beyond simple acceptance prediction, researchers began exploring automatic review generation. Bartoli et al. (2016) proposed one of the earliest neural frameworks, trained on 48 papers from their lab, to generate review comments. Nagata (2019) generated sentence-level feedback on grammar errors to support academic writing. ReviewRobot (Wang et al. 2020) constructed a knowledge graph from the input paper, predicted review scores, and selected templated comments based on both scores and supporting evidence. However, traditional neural models struggled with long and technical documents, lacking the capacity for deep semantic understanding, which ultimately hindered progress in this field.
@@
Recent advancements in LLMs have inspired a growing body of work on LLM-based automatic reviewing. Multi-agent collaboration frameworks aim to simulate the multi-role dynamics of real-world peer review. For example, ReviewMT (Tan et al. 2024) and AIScientist (Lu et al. 2024) reframe the review process as a multi-round, long-context dialogue among multiple roles, i.e., reviewers, authors, and meta-reviewers. AgentReview (Jin et al. 2024) further explores this direction by prompting LLMs to assume diverse reviewer personalities and decision strategies. Fine-tuned reviewer models enhance alignment with human review standards through supervised training. CycleReviewer (Weng et al. 2025) fine-tunes an open-source LLM on domain-specific review data, simulating multiple reviewers who assess the paper across different dimensions. DeepReview (Zhu et al. 2025) extends this line of work by modeling multi-step reviewer reasoning and training on generated rationales to produce more coherent, logically grounded reviews. Graph-based approaches such as GraphEval (Feng et al. 2025) leverage LLM prompting to segment abstracts into discrete opinion sentences, which are then connected via similarity-based edges and processed through graph reasoning to predict acceptance decisions.
@@
Figure 1: The workflow of ReViewGraph. Reviewer1, Reviewer2, and Reviewer3 represent three regular reviewer agents. Nodes are typed by shape: circles denote paper topic nodes, pentagrams denote evaluation dimension nodes (e.g., Methodological Novelty), squares denote opinions of the reviewers, and triangles represent the author’s opinions.
@@
To represent the complex reasoning and interaction dynamics of the peer review process, we construct a heterogeneous debate graph from simulated reviewer-author interactions. Following the standard formulation of heterogeneous information networks (Hu et al. 2020), we define the graph as $G=\{\mathcal{V},\mathcal{E},\mathcal{A},\mathcal{R}\}$, where $\mathcal{V}$ is the set of nodes, $\mathcal{E}$ is the set of edges, $\mathcal{A}$ is the set of node types, and $\mathcal{R}$ is the set of edge types. Each node $v\in\mathcal{V}$ and $e\in\mathcal{E}$ is associated with a type via node and edge type mapping functions: $\psi(v):V\rightarrow\mathcal{A}$, $\eta(e):E\rightarrow\mathcal{R}$. Each edge corresponds to a meta-relation in the form $\psi(s),\eta(e),\psi(t)>$, where $s$ and $t$ are source and target nodes.
@@
Following prior works (Jin et al. 2024; Zhu et al. 2025; Weng et al. 2025; Feng et al. 2025), we use the full paper as input and aim to predict the final review decision. The decision $s$ is selected from a predefined set of review outcomes: accept or reject.
@@
We compare our proposed framework against several representative baselines that leverage LLMs for automatic paper review: (1) ICL-based Method (Brown et al. 2020): Directly prompts a pre-trained LLM using in-context examples to predict the review decision based on the full paper content. (2) CoT-based Method (Wei et al. 2022): Encourages the LLM to perform chain-of-thought reasoning before generating the final review outcome. (3) AI-Scientist (Lu et al. 2024): Redefines the review process as a multi-turn, long-context dialogue among multiple roles. (4) CycleReviewer (Weng et al. 2025): Fine-tunes an open-source LLM using domain-specific peer review data. (5) DeepReview (Zhu et al. 2025): Extends CycleReviewer by modeling multi-step reviewer reasoning and training on generated rationales. (6) GraphEval (Feng et al. 2025): Segments paper abstracts into discrete opinion sentences using LLM prompting, connects them via similarity-based edges to construct a graph, and applies graph-based reasoning to predict the final review decision. The implementation details of ReViewGraph and the baselines are provided in the source code repository.
@@
Case Study.
To analyze the advantages of our method, we present two representative cases as follows: Case 1: Correct Rejection Prediction Amid Subtle Negative Consensus. In the first case shown in Figure 2(a), the ground-truth decision was Reject. While the paper proposed a novel approach, multiple reviewers expressed concerns about the lack of rigorous theoretical analysis and inadequate comparisons with state-of-the-art baselines. Reviewer 1 noted the absence of evidence demonstrating the solution’s diversity. Reviewer 2 emphasized the theoretical gap in stochastic settings and the lack of comparison with methods designed for the $n\gg m$ scenario. Reviewer 3 echoed similar concerns regarding limited convergence analysis and experimental comparisons. Although the authors acknowledged these issues and promised improvements, no concrete resolutions were provided. ReViewGraph successfully captured the negative consensus across multiple evaluation dimensions (e.g., theoretical rigor, baseline comparison) by modeling reviewer–reviewer agreement and reviewer–author interactions in a structured heterogeneous graph. The system correctly identified the dominant negative stance despite superficially polite language or minor positive comments, and inferred the rejection outcome. In contrast, methods such as ICL, CoT prompting, and LLM-based dialogue agents misclassified the case as Accept, likely due to over-reliance on isolated affirmative signals. Case 2: Correct Acceptance Prediction Despite Isolated Criticism. In another instance shown in Figure 2(b), the ground-truth decision was Accept. Reviewer 2 suggested further exploration of the impact of calibration techniques on computational efficiency. However, both Reviewer 1 and Reviewer 3 explicitly stated that the paper adequately addressed computational concerns and praised its theoretical guarantees. Reviewer 2 ultimately agreed with the decision to accept. The authors also acknowledged the suggestion and expressed willingness to expand the analysis in future revisions. While several baseline models incorrectly predicted a Reject decision, possibly misinterpreting Reviewer 2’s suggestion as a critical flaw, our model accurately inferred that Reviewer 2’s comment was an isolated and non-decisive concern. By modeling the majority–minority stance across reviewers and the consistency between evaluation dimensions and final decisions, our method avoided overweighing minority dissent.
@@ Conclusion
We proposed ReViewGraph, a novel framework for automatic paper reviewing that leveraged heterogeneous graphs constructed from simulated reviewer–author interactions. By modeling multi-agent debates, we extracted structured opinion relations among reviewers and authors and represented them as a heterogeneous graph with semantically typed nodes and edges. A heterogeneous graph Transformer was then applied to reason over this structure and predict final review outcomes. Through comprehensive experiments on three ICLR datasets, ReViewGraph consistently outperformed strong prompt-based, fine-tuned, and graph-based baselines across multiple evaluation metrics. Notably, our method achieved a relative improvement of 15.73% in Macro F1 score over the best-performing baseline, while requiring no LLM fine-tuning. Ablation studies further validated the importance of each graph component, particularly the modeling of evaluation dimensions and inter-agent relations. Case studies also demonstrated that ReViewGraph was able to correctly interpret nuanced reviewer dynamics, such as subtle consensus or isolated dissent, which were often misclassified by other methods. These findings highlighted that explicitly modeling reviewer–author interactions and discourse-level semantics led to more robust, interpretable, and context-aware automated reviewing. Overall, ReViewGraph offered a scalable and effective framework for augmenting peer review with structured LLM reasoning, pointing toward future directions in trustworthy AI-assisted scientific evaluation.
@@ Acknowledgments
We would like to thank the anonymous reviewers for their valuable comments and suggestions to improve this paper. Min Yang was supported by National Key Research and Development Program of China (2024YFF0908201), National Natural Science Foundation of China (Grant No. 62376262), the Natural Science Foundation of Guangdong Province of China (2024A1515030166, 2025B1515020032). Xian Wei is supported by the General Program of Shanghai Natural Science Foundation (No.24ZR1419800, No.23ZR1419300), the National Natural Science Foundation of China (No.42130112, No.42371479), the Science and Technology Commission of Shanghai Municipality (No.22DZ2229004), and Shanghai Frontiers Science Center of Molecule Intelligent Syntheses. Shiwen Ni was supported by GuangDong Basic and Applied Basic Research Foundation (2023A1515110718 and 2024A1515012003), China Postdoctoral Science Foundation (2024M753398), Postdoctoral Fellowship Program of CPSF (GZC20232873) and Shenzhen Science and Technology Program (JCYJ20250604182917023).
@@
For structured reasoning over reviewer-author interactions, we design a heterogeneous graph and apply a two-layer Heterogeneous Graph Transformer (HGT) with a total of three layers. Each layer uses multi-head attention, and the hidden dimension is set to 128. To initialize node and edge representations, we use Qwen3-Embedding [^6] https://huggingface.co/Qwen/Qwen3-Embedding-8B. The model is trained for up to 100 epochs with a batch size of 32, using early stopping based on validation F1 score. Optimization is performed using the Adam optimizer (Kingma and Ba 2015) with a learning rate of $1\times 10^{-4}$. Our code and data are provided in the supplementary materials.
@@
Table 9 and Table 10 respectively summarize the prompt construction and implementation details of the ICL-based (Brown et al. 2020) and CoT-based methods (Wei et al. 2022). Table 11 and Table 12 present the prompt templates used by the AI-Scientist (Lu et al. 2024) method for paper reviewing and review aggregation, respectively. All baselines are run using consistent inference settings (e.g., temperature, max tokens) unless otherwise stated.
For the GraphEval (Feng et al. 2025) baseline, we follow the original implementation to preprocess our data and apply the same model configuration for training and evaluation [^7] https://github.com/ulab-uiuc/GraphEval. Specifically, we adopt a Graph Convolutional Network (GCN) as the backbone, configured as a two-layer weighted GNN with a hidden dimension of 64. The model is trained for up to 100 epochs with a batch size of 32, using the Adam optimizer.
For DeepReview (Zhu et al. 2025) and CycleReview (Weng et al. 2025), we directly invoke the publicly released models, DeepReviewer-14B [^8] https://huggingface.co/WestlakeNLP/DeepReviewer-14B, CycleReviewer-ML-LLaMA-3.1-8B [^9] https://huggingface.co/WestlakeNLP/CycleReviewer-ML-Llama-3.1-8B, and CycleReviewer-LLaMA-3.1-70B [^10] https://huggingface.co/WestlakeNLP/CycleReviewer-Llama-3.1-70B, without any fine-tuning or additional adaptation. This ensures a fair and reproducible baseline comparison, highlighting the benefits of our explicitly structured multi-agent debate framework.